Let M = 8.00kg, m = 6.00kg, θ = 40.00, and the coefficient of kinetic friction between m and the incline plane is 0.40 ; then after releasing the system what is the acceleration of each object? The pulley and string are small.
Answer #1

From you question statement, I assume that mass M is hanging and mass m is on an incline....right ????

Let me know if diagram is different.

From free body diagram of M,

Mg = T

From free body diagram of m,

mgsin - ukmgcos = ma

Therefore, total net force

F = Mg - (mgsin + ukmgcos)

F = 8*9.8 - (6*9.8*sin 40 + 0.4*6*9.8*cos 40)

F = 22.586 N

so,

a = F / total mass

a = 22.586 / 14

a = 1.613 m/s2